# DESI Dark Energy Time Evolution is Recovered by Cosmologically Coupled Black Holes

### v1

@@ Abstract
Recent results from baryon acoustic oscillation measurements by the Dark Energy Spectroscopic Instrument (DESI) have found preliminary evidence that dark energy (DE) evolves with time, as parameterized by a $w_{0}w_{a}$ equation of state. In this study, we point out that the DESI-preferred $w_{0}w_{a}$ models are recovered when a $w_{0}w_{a}$ model is fit to DE produced by baryon conversion in cosmologically coupled black holes (BHs). This recovery does not require any ad hoc parameter adjustments; it solely depends on the independently measured cosmic star formation rate density. We discuss our result in the context of the missing baryon problem and the anomalously low sum of neutrino masses preferred by DESI. The global evolution of DE is an orthogonal probe of cosmological coupling, complementing constraints on BH mass growth from elliptical galaxies, stellar binaries, globular clusters, the LIGO-Virgo-KAGRA merging population, and X-ray binaries. A DE density that correlates with star formation is a natural outcome of cosmological coupling in BH populations.
@@ 1 Introduction
Recently, the Dark Energy Spectroscopic Instrument (DESI), the first Stage IV spectroscopic survey, reported significantly enhanced cosmology constraints from Baryon Acoustic Oscillation (BAO) measurements using its first year of data (DESI-1YR) [8]. The DESI collaboration finds that adopting a $w_{0}w_{a}$ DE equation of state and combining DESI-1YR BAO data with CMB lensing data and Type Ia SNe data from the Pantheon+, Union3 and DES-5YR datasets leads to deviations from $\Lambda$CDM by $2.5\sigma$, $3.5\sigma$ and $3.9\sigma$, respectively. The values of $w_{0}=-0.831\pm{0.066},-0.64\pm{0.011}$ and $-0.724\pm{0.071}$ and $w_{a}=-0.73^{+0.32}_{-0.28},-1.30^{+0.45}_{-0.39}$ and $-1.06^{+0.35}_{-0.31}$ for the combinations DESI + CMB + Pantheon+, Union3 and DES-5YR respectively, indicate a strong preference for $w_{0}>-1$ and large $w_{a}<0$. Substituting these values into the $w_{0}w_{a}$ parameterization, Eq. 1.1 implies that, at some point in the past, $w<-1$.
@@
An equation of state $w<-1$ is often referred to as “phantom,” because it appears to generate energy densities from nothing. [^1] Phantom models typically require a negative kinetic energy term, which can be envisioned as a scalar field rolling up its potential [9]. In addition, phantom models have been argued to have theoretical inconsistencies [10, 11, 12, 13]. Because this would imply an unstable vacuum, the interpretation of DESI data within the context of the $\Lambda$CDM model [14, 15, 16, 17, 18, 19], as well as suggestions for interacting DE [20], quintessence scalar fields [21, 22, 23], and other scenarios beyond $\Lambda$CDM [24, 25], has attracted recent interest from the scientific community. Phantom DE is also considered a useful phenomenological description of models in which DE actually receives energy from another species. Models have been proposed where dark matter (DM) couples to DE [26, 27, 28, 29, 30, 31]. However, very little is known about DM, and it has yet to be detected directly [32, 33, 34, 35, 36], which currently limits the predictive power of these models for the phenomenology of DE.
@@
Another possibility involves models where baryons couple to DE. For example, the conversion of DE into baryons explains the origin of the Big Bang in the inflationary paradigm [37, 38]. Conversely, the conversion of baryons into DE was proposed by Gliner to occur during the gravitational collapse of dead stars, as described in the paper that theoretically discovered what we now call DE [39]. Following this study, a diverse community of researchers has developed exact GR solutions of BHs with DE interiors [40, 41, 42, 43, 44] and explored condensed-matter and quantum field-theoretic interpretations (see e.g. [45, 46, 47]). These BH solutions often mimic the familiar Schwarzschild and Kerr BH models but are singularity-free and do not have horizons. This can lead to surprising phenomena when these objects are generalized to cosmological boundary conditions. For instance, there are known GR solutions for BHs [e.g. 48, 49] that expand in lockstep with their embedding cosmology, gaining mass independently of accretion or merger. [^2] Recent theoretical objections to cosmological coupling [50, 51, 52, 53, 54] demonstrate the existence of decoupled solutions. Because GR is a non-linear theory, the existence of a solution does not imply uniqueness, (see e.g. §1.2 of [55]) and so cannot exclude other solutions. Solutions with realistic cosmological boundary conditions and mass growth $\propto a$ establish these phenomena as a robust GR prediction [49, 56]. Coupling also follows from a mathematically rigorous treatment of the Einstein-Hilbert action for perturbed Robertson-Walker spacetime, where metric degrees of freedom are necessarily constrained in Fourier space [57, 58, 59].
@@
Recently, observational evidence for purely cosmological mass growth has been reported in the super-massive black holes (SMBHs) of elliptical galaxies [60]. The preferred rate of mass growth for this population, $m\propto a^{3}$ [61] is consistent with observations from NANOgrav [62] and low-mass X-ray binaries (XRBs) [63]. Because cosmological number densities decrease $\propto 1/a^{3}$, if SMBHs gain mass $\propto a^{3}$, then (in the absence of production) their physical density is constant. Conservation of stress energy then requires that they enter Friedmann’s equations as DE, consistent with the DE-interior BH models. However, BH growth $\propto a^{3}$ exceeds the upper bounds $\propto a$, from the known coupled GR solutions and recent measurements in stellar-mass BH populations of LIGO-Virgo-KAGRA [64, 65, 66, 67], globular clusters [68], Gaia binaries [69], and high-mass XRBs [70]. Local measurements of cosmological mass growth are complicated by accretion and merger, and require accurate estimation of BH formation time, making orthogonal probes of this GR effect valuable for understanding the emerging theoretical and observational landscape.
@@
DESI and forthcoming Stage IV surveys [71, 72, 73, 74, 75, 76, 77, 78] can provide positive evidence for DE within cosmologically coupled BHs. The cosmological background offers an averaged dynamics, free from selection effects and assumptions about minimum BH mass [^3] For BH models without horizons, the Tolman-Oppenheimer-Volkoff stability limit $\lesssim 2.2M_{\odot}$ for maximal neutron star mass can no longer be used to establish a lower-bound BH mass. This also confounds interpretation of low-mass remnants without optical or merger tidal-deformation signatures, e.g. [79]. required to constrain coupled mass growth in isolated stellar binaries. If BHs contribute as a DE species, then DE time-evolution will track BH production and growth. [^4] The possible role of stellar-collapse BHs in cosmic expansion was first proposed in the Gravitational Aether Theory, an extension to GR [80, 81]. Because stellar-collapse is expected to be the dominant BH formation channel over the DESI-sensitive redshift range $z\lesssim 3$, and this range overlaps with the peak rate of cosmic star formation around $z\sim 2$, DESI is well-positioned to search for a cosmologically coupled BH population. In this paper, we present the first search for this signal at “cosmic noon.” Because the unprocessed DESI-1YR data has not yet been publicly released, we analyze DE as sourced by cosmologically coupled stellar-collapse BHs (BH DE) within the $w_{0}w_{a}$ framework. We find that the $w_{0}w_{a}$ evolution expected from BH DE matches the DESI most-likely $w_{0}w_{a}$ models when those models include any SNe datasets.
Figure 1: Agreement between the $w_{0}w_{a}$ best-fit to the BH DE model and all combinations of DESI datasets. (Dotted) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ prediction of the BH DE model, where $w:=-1$ and DE is sourced by non-singular stellar collapse BHs, as described in Eq. 2.11. DESI uses a $w_{0}w_{a}$ parameterization, which limits its ability to fit predictions of the BH DE model. (Dashed) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ prediction given by the least-squares best-fit $w_{0}w_{a}$ parameterization to the BH DE model. (Solid) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ implied by DESI measurements of $w_{0}$ and $w_{a}$. Each panel is determined from the DESI-1YR central values of $w_{0}$, $w_{a}$, $\Omega_{m}$, and $H_{0}$ for DESI + CMB + Pantheon+ (top), DESI + CMB + Union3 (middle), and DESI + CMB + DESY5 (bottom) (see Table 3 in [82]). In all panels, the best-fit DESI model closely tracks the least-squares best-fit to the BH DE model with a standard star formation rate density [83]. Therefore, DESI observations are consistent with $w=-1$, provided that DE is produced by stellar-collapse to non-singular, cosmologically coupled, BHs.
@@ 2 Theory
We adopt the position that BHs are non-singular vacuum energy objects, cosmologically coupled, and produced solely by stellar collapse. To describe this process phenomenologically, we define the physical baryon density as follows:
$\rho_{b}:=\begin{cases}\displaystyle\frac{\rho_{\mathrm{cr}}\Omega_{b}}{a^{3}}&a<a_{i}\\ \displaystyle\frac{\rho_{\mathrm{cr}}\Omega_{b}}{a^{3}}-\frac{\Xi}{a^{3}}\int_{a_{i}}^{a}\psi\frac{\mathrm{d}a^{\prime}}{Ha^{\prime}}&a\geqslant a_{i}\end{cases}\,.$ (2.1)
Here, $\psi$ represents a model for the comoving star formation rate density (SFRD). The parameters $\rho_{\mathrm{cr}}$ and $\Omega_{b}$ denote the critical density and baryon fraction, respectively, inferred from early-universe measurements projected to the present day. In essence, the physical baryon density evolves as inferred by, for example, Planck, until the onset of stellar first light at $a_{i}$. After this point, a constant fraction $\Xi$ of baryons are depleted in synchrony with the assembly of stellar mass. The assumption that $\Xi$ is constant reflects the simplification that its time evolution is subdominant to the time evolution of $\psi$. We generically define the DE contribution through its pressure as:
$P_{\Lambda}:=w\rho_{\Lambda}\,,$ (2.2)

### v2

@@ Abstract
Recent baryon acoustic oscillation (BAO) measurements by the Dark Energy Spectroscopic Instrument (DESI) provide evidence that dark energy (DE) evolves with time, as parameterized by a $w_{0}w_{a}$ equation of state. Cosmologically coupled black holes (BHs) provide a DE source that naturally evolves with time, because BH production tracks cosmic star-formation. Using DESI BAO measurements and priors informed by Big Bang Nucleosynthesis, we measure the fraction of baryonic density converted into BHs, assuming that all DE is sourced by BH production. We find that the best-fit DE density tracks each DESI best-fit $w_{0}w_{a}$ model within $1\sigma$, except at redshifts $z\lesssim 0.2$, highlighting limitations of the $w_{0}w_{a}$ parameterization. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)~{}\,{\rm km\,s^{-1}\,Mpc^{-1}}$, in excellent agreement with the Chicago-Carnegie Hubble Program, with the same $\chi^{2}$ as $\Lambda$CDM, and with two fewer parameters than $w_{0}w_{a}$. Because cosmologically coupled BH production depletes the baryon density established by primordial nucleosynthesis, these BHs provide a physical explanation for the “missing baryon problem” and the anomalously low sum of neutrino masses preferred by DESI. The global evolution of DE is an orthogonal probe of cosmological coupling, complementing constraints on BH mass-growth from elliptical galaxies, stellar binaries, globular clusters, the LIGO-Virgo-KAGRA merging population, and X-ray binaries. A DE density that correlates with cosmic star-formation: 1) is a natural outcome of cosmological coupling in BH populations; 2) eases tension between early and late-time cosmological probes; and 3) produces time-evolution toward a late-time $\Lambda$CDM cosmology different from Cosmic Microwave Background projections.
@@ 1 Introduction
Figure 1: Dark energy (DE) density as a function of redshift predicted from cosmologically coupled, stellar-collapse BHs (black). Because stellar collapse BH production tracks star-formation, the impact of two empirically determined models is shown. The Trinca $\psi$ model (solid) features ample star-formation at high $z$, consistent with recent James Webb Space Telescope observations [4, 5]. The Madau $\psi$ model (dashed) is a fiducial rate determined from numerous IR and UV luminosity measurements primarily from $z<4$ [6, 7]. DESI parameterizes DE time-evolution via equation of state parameters $w_{0}$ and $w_{a}$, as determined from joint fits to BAO, CMB, and SNe datasets. Regions (shaded) reflect ±$\pm 1\sigma$ uncertainties in DE density, as determined by SNe datasets from Pantheon (green), Union3 (red), and DES Y5 (blue). The time-evolution of DE density, when sourced by stellar-collapse BHs, remains within $<1\sigma$ of the $w_{0}w_{a}$ model for each dataset until $z\lesssim 0.2$. No combination of $w_{0}$ and $w_{a}$ can lead to a DE density that approaches a constant value from below. This asymptoting behavior of late-time DE is expected from cosmologically coupled BHs, because the production of stellar progenitors has been decreasing for the past $\sim 9$ Gyr.
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI), the first Stage IV spectroscopic survey, reported significantly enhanced cosmology constraints from Baryon Acoustic Oscillation (BAO) measurements using its first year of data [12]. The DESI collaboration finds that adopting a $w_{0}w_{a}$ DE equation of state and combining: DESI-1YR BAO data; Cosmic Microwave Background (CMB) lensing data [13]; and Type Ia SNe data from the Pantheon+ [11], Union3[9] and DES-5YR [10] datasets leads to deviations from $\Lambda$CDM by $2.5\sigma$, $3.5\sigma$ and $3.9\sigma$, respectively, indicating a strong preference for $w_{0}>-1$ and large $w_{a}<0$. These results can be understood by examining the DE density implied by Eq. 1.1,
$\rho_{\mathrm{DE}}\propto\frac{e^{3aw_{a}}}{a^{3(1+w_{0}+w_{a})}}\,.$ (1.2)
@@
Note that if $01+w_{0}+w_{a}>$ 0, then the denominator would diverge in the early universe as $a\to 0$. Thus $1+w_{0}+w_{a}\leqslant 0$ must always be true, and the DE density initially increases from zero at $a=0$. The exponential in the numerator then guarantees a “hump” for $w_{a}<0$. Other physical behaviors, like “asymptotes” and “multiple humps,” cannot be described by Eq. 1.1. The DESI data may already highlight the provisional nature of Eq. 1.1, because the preferred values for $w_{0}$ and $w_{a}$ allow causality violation and generate energy densities from nothing. [^1] An equation of state $w<-1$ is often referred to as “phantom” for this reason. Phantom models typically require a negative kinetic energy term, which can be envisioned as a scalar field rolling up its potential [14]. In addition, phantom models have been argued to have theoretical inconsistencies [15, 16, 17, 18]. Because this would imply an unstable vacuum, the interpretation of DESI data within the context of the $\Lambda$CDM model [19, 20, 21, 22, 23, 24], as well as suggestions for interacting DE [25], quintessence scalar fields [26, 27, 28], and other scenarios beyond $\Lambda$CDM [29, 30], has attracted recent interest from the scientific community.
@@
The community has established that both of these theoretical obstacles can be removed if DE actually does receive energy from another species. Models have been proposed where dark matter (DM) couples to DE [31, 32, 33, 34, 35, 36, 37]. Very little is known about DM, however, and it has yet to be detected directly [38, 39, 40, 41, 42], which currently limits the predictive power of these models for the phenomenology of DE. Another possibility involves models where baryons couple to DE. For example, the conversion of DE into baryons explains the origin of the Big Bang in the inflationary paradigm [43, 44]. The reverse process, the conversion of baryons into DE, was proposed by Gliner to occur during the gravitational collapse of dead stars in the paper that theoretically anticipated DE’s discovery [45]. Following this study, a diverse community of researchers has developed exact GR solutions of BHs with DE interiors [46, 47, 48, 49, 50] and explored condensed-matter and quantum field-theoretic interpretations (see e.g. [51, 52, 53]). These BH solutions often mimic the familiar Schwarzschild and Kerr BH models for times $\ll$ the Hubble time, but are singularity-free and need not have horizons. This can lead to surprising phenomena when these objects are generalized to astrophysically realistic boundary conditions. For instance, there are known GR solutions for BHs [e.g. 54, 55] that expand in lockstep with their embedding cosmology, gaining mass independently of accretion or merger. [^2] Recent theoretical objections to cosmological coupling [56, 57, 58, 59, 60] demonstrate the existence of decoupled solutions. Because GR is a non-linear theory, the existence of a solution does not imply uniqueness, (see e.g. §1.2 of [61]) and so cannot exclude other solutions. Solutions with realistic cosmological boundary conditions and mass growth $\propto a$ establish these phenomena as a robust GR prediction [55, 62]. Coupling also follows from a mathematically rigorous treatment of the Einstein-Hilbert action for perturbed Robertson-Walker spacetime, where metric degrees of freedom are necessarily constrained in Fourier space [63, 64, 65]. It has also been shown that the boundary of a BH must couple to any time-dependent exterior geometry, or else result in a naked null singularity [66]. This result is independent of the particular field equations, e.g. Einstein’s, and applies to any (single) metric theory.
@@
Recently, observational evidence for cosmological mass growth has been reported in the super-massive black holes (SMBHs) of massive early-type galaxies[67, 67]. The preferred rate of mass growth for this population, $m\propto a^{3}$ [68] is consistent with the local SMBH mass density proposed by NANOgrav [69] and low-mass X-ray binaries (XRBs) [70]. Because cosmological number densities decrease $\propto 1/a^{3}$, if all BHs gain mass $\propto a^{3}$, then (in the absence of production) their physical density is constant. Conservation of stress energy then requires that they enter Friedmann’s equations as DE (i.e. with pressure $P=-\rho$), consistent with the DE-interior BH models. However, BH growth $\propto a^{3}$ exceeds upper bounds $\propto a$, determined from known coupled GR solutions and recent measurements in stellar-mass BH populations of LIGO-Virgo-KAGRA [71, 72, 73, 74], globular clusters [75], Gaia binaries [76], and high-mass XRBs [77]. Local measurements of cosmological mass growth are complicated by uncertainties in accretion, merger history, and formation time, making orthogonal probes of this GR effect valuable for understanding the emerging theoretical and observational landscape.
@@
DESI and forthcoming Stage IV surveys [78, 79, 80, 81, 82, 83, 84, 85] can provide positive evidence for DE within cosmologically coupled BHs (CCBHs). The cosmological background offers an averaged dynamics, free from selection effects and assumptions about minimum BH mass [^3] For BH models without horizons, the Tolman-Oppenheimer-Volkoff stability limit $\lesssim 2.2M_{\odot}$ for maximal neutron star mass can no longer be used to establish a lower-bound BH mass. This also confounds interpretation of low-mass remnants without optical or merger tidal-deformation signatures, e.g. [86]. required to constrain coupled mass growth in isolated stellar binaries. If BHs contribute as a DE species, then DE time-evolution will track BH production and growth. [^4] The possible role of stellar-collapse BHs in cosmic expansion was first proposed in the Gravitational Aether Theory, an extension to GR [87, 88]. In particular, star-formation peaks at $z\sim 2$ [6, 7] and then drops by a factor of 10$\sim$ 10$\times$ 10 × by today. This causes the DE density from CCBHs to asymptote toward a constant value at $z\lesssim 1$. This would mimic $\Lambda$CDM in the late-time universe, but with a DE density and expansion history that disagrees with CMB projections. Because stellar-collapse is expected to be the dominant BH formation channel over the DESI-sensitive redshift range $z\lesssim 3$, and this range overlaps with the peak rate of cosmic star-formation, DESI is well-positioned to distinguish a CCBH population from other hypotheses.
In this paper, we present the first search for significant DE production at “cosmic noon.” In Section 2 we implement baryon depletion, which tracks the cosmic star-formation rate within Friedmann cosmology. In Section 3, we detail our methodology for likelihood and parameter estimation with the DESI BAO measurements. In Section 4, we present our main results. In Section 5, we discuss implications of cosmologically coupled DE for the current census of baryons in the local universe, the anomalously low summed neutrino mass preferred by DESI, and currently understood BH and accretion physics. Because we will be describing and comparing different model fit to different datasets, we will consistently refer to these combinations with “Model (Datasets)” notation.
@@ 2 Theory
We adopt the position that BHs are non-singular vacuum energy objects, cosmologically coupled, and produced solely by stellar collapse [c.f. 89]. Because the DE density is determined dynamically, defining the cosmology with the density parameters $\Omega_{i}$ and a distinct present-day Hubble rate $H_{0}$ is no longer straightforward. To keep as close to standard notation and usage as possible, we instead focus on “little $\omega$’s,” which are typically defined through $\omega_{b}:=\Omega_{b}h^{2}$. This definition implies that,
$\rho_{i}\bigg{|}_{a=1}=\rho_{\mathrm{cr}}\Omega_{i}=\left(\frac{3H_{0}^{2}}{8\pi G}\right)\Omega_{i}=\left(\frac{3\times 10^{4}}{8\pi G}\right)\omega_{i}:=C\omega_{i}.$ (2.1)
For matter species, the $C\omega_{i}$ are comoving energy densities. In contrast to $\Omega_{i}$ and $H_{0}$, they are directly measured by early-universe probes: $\omega_{c}$ and $\omega_{b}$ are the two background parameters of $\Lambda$CDM. In what follows, we will define $C\omega_{b}^{\mathrm{proj}}$ to be the projected comoving density of baryons, as inferred from early-universe measurements. We may now define baryon consumption phenomenologically via the physical baryon density:
$\rho_{b}:=\begin{cases}\displaystyle\frac{C\omega_{b}^{\mathrm{proj}}}{a^{3}}&a<a_{i}\\ \displaystyle\frac{C\omega_{b}^{\mathrm{proj}}}{a^{3}}-\frac{\Xi}{a^{3}}\int_{a_{i}}^{a}\psi\frac{\mathrm{d}a^{\prime}}{Ha^{\prime}}&a\geqslant a_{i}\end{cases}\,.$ (2.2)
Here, $\psi$ is the observed comoving star-formation rate density (SFRD), with units comoving density per unit proper time, consistent with $C\omega_{b}^{\mathrm{proj}}$. Instead, the physical baryon density $\rho_{b}$ evolves $\propto\omega_{b}^{\mathrm{proj}}/a^{3}$, as inferred by Planck, until the onset of stellar first light at $a_{i}$. Hereafter, we define $\Xi$ as the fraction of baryons depleted in synchrony with the assembly of stellar mass. The assumption that $\Xi$ is constant reflects the simplification that its time evolution is subdominant to the time evolution of $\psi$.
We generically define the DE contribution through its pressure as:
$P_{\mathrm{DE}}:=w\rho_{\mathrm{DE}}\,,$ (2.3)